# Model Informed Flows for Bayesian Inference of Probabilistic Programs

### v1

@@ Abstract
Variational inference often struggles with the posterior geometry exhibited by complex hierarchical Bayesian models. Recent advances in flow-based variational families and Variationally Inferred Parameters (VIP) each address aspects of this challenge, but their formal relationship is unexplored. Here, we prove that the combination of VIP and a full-rank Gaussian can be represented exactly as a forward autoregressive flow augmented with a translation term and input from the model’s prior. Guided by this theoretical insight, we introduce the Model-Informed Flow (MIF) architecture, which adds the necessary translation mechanism, prior information, and hierarchical ordering. Empirically, MIF delivers tighter posterior approximations and matches or exceeds state-of-the-art performance across a suite of hierarchical and non-hierarchical benchmarks.
@@ 1 Introduction
Inference remains challenging for many complex Bayesian models. Variational Inference (VI) casts posterior approximation as minimization of the Kullback-Leibler divergence between a tractable approximating distribution and the true posterior (blei2017variational, ; jordan1999introduction, ; hoffman2013stochastic, ). Recent advances have focused on increasing the expressiveness of variational families. Flow-based models, which apply a sequence of invertible transformations to a simple base distribution, have shown particular promise (papamakarios2017masked, ; rezende2015variational, ; kingma2016improved, ).
Variationally Inferred Parameters (VIP) (gorinova2020automatic, ) address a key challenge in VI: The posteriors of hierarchical Bayesian models often exhibit pronounced curvature or “funnel-like” geometry (betancourt2015hamiltonian, ; ingraham2017variational, ) that standard families like Gaussians cannot capture. VIP builds on insights from non-centered parameterization (NCP) in Markov chain Monte Carlo (MCMC) (papaspiliopoulos2007general, ), adaptively learning the optimal degree of non-centering for each latent variable during the VI optimization process.
Our preliminary investigations reveal a striking observation: When VIP is combined with full-rank Gaussian variational families, it achieves performance comparable to state-of-the-art methods from recent evaluation (blessing2024beyond, ) (See Tables 1 and  3). Since VIP and flow-based models both excel at navigating posterior geometry, might they be manifestations of a common principle? In particular, can a flow-based distribution encapsulate the benefits of VIP?
@@ 2.2 Hierarchical Bayesian Models
Hierarchical Bayesian models are common in domains such as ecology (cressie2009accounting, ; wikle2003hierarchical, ; royle2008hierarchical, ) and epidemiology (dominici2000combining, ; richardson2003bayesian, ; lawson2018bayesian, ). Following Gorinova et al. gorinova2020automatic , we consider a hierarchical model in which each latent variable $z_{i}$ is conditionally Gaussian with a mean and variance that are arbitrary function of parent variables, while the observed data $x_{j}$ is generated from an arbitrary likelihood $p(x_{j}|\pi(x_{j}))$. Formally, write
@@
where $\pi(\cdot)$ denotes the set of parent variables. In practice, the functions $f_{i}$ and $g_{i}$ (which return the mean and standard deviation of $z_{i}$, respectively) often arise directly from a probabilistic programming system, encoding how the mean and variance of each latent variable depend on its parent variables.
@@ 2.3 Non-Centered Parameterization
Hierarchical models often exhibit difficult curvature and “funnel-like“ geometries. These are a challenge for both sampling and variational inference approaches (betancourt2015hamiltonian, ; ingraham2017variational, ; hoffman2019neutra, ). A common strategy in MCMC to alleviate these issues is the non-centered parameterization (NCP) (papaspiliopoulos2007general, ). To illustrate, consider two latent variables $z_{1}$ and $z_{2}$ having prior of
@@
In the standard (centered) parameterization, the goal is to approximate $p(z|x)$. However, the nonlinear dependence of $z_{2}$ on $z_{1}$ can yield a sharply funnel-shaped joint posterior. As illustrated in Figure 1 (left panel), a full-rank Gaussian struggles to capture this funnel structure.
@@ 2.4 Variationally Inferred Parameters
then recovers $z_{i}$ via VIP transformation, $T_{\mathrm{VIP}}$, as in Definition 1. Note that $\lambda_{i}$ interpolates between the centered ($\lambda_{i}=1$) and fully non-centered ($\lambda_{i}=0$) extremes.
@@
For a fixed ordering of latent variables $z=(z_{1},\dots,z_{D})$, and auxiliary variables $\tilde{z}=(\tilde{z}_{1},\dots,\tilde{z}_{D})$, let $\lambda_{i}\in[0,1]$ be the partial non-centering parameter for each coordinate $i$, and $f_{i}(\pi(z_{i}))$ and $g_{i}(\pi(z_{i}))$ be continuous functions that return the mean and standard deviation of $z_{i}$ (respectively) based on the parent nodes $\pi(z_{i})$. Then, the VIP transformation $z=T_{\mathrm{VIP}}(\tilde{z})$ is defined coordinate-wise by
@@
Substituting $T_{\mathrm{VIP}}$ into the joint distribution (Equation 2.3) yields a partially non-centered representation
@@
VIP then learns a variational distribution $q_{w}(\tilde{z})$ by minimizing $\mathrm{KL}\left(q_{w}(\tilde{z})\|p_{\mathrm{VIP}}(\tilde{z}\mid x)\right)$, optimizing both the variational parameters $w$ and the partiality parameters $\lambda$.
@@ 2.5 Equivalence of VIP in Model vs Variational Spaces
VIP applies $T_{\mathrm{VIP}}$, in the model space. However, there is a straightforward analog in the variational space. Suppose $q_{w}(\tilde{z})$ is any distribution over the non-centered latent variables $\tilde{z}$. We define a new distribution $q_{w,\mathrm{VIP}}(z)$ by letting
$z=T_{\mathrm{VIP}}(\tilde{z},x)\quad\text{ with }\quad\tilde{z}\sim q_{w}(\tilde{z}).$ (2.9)
@@
Since the KL divergence is invariant under such invertible transformations (cover1991elements, ), we have that
@@
Thus, optimizing $\min_{w,\lambda}\mathrm{KL}\left(q_{w,\mathrm{VIP}}(z)\|p(z|x)\right)$ with the posterior fixed is equivalent to reparameterizing the posterior as in 2.4. In this paper, we choose a Gaussian distribution for $q_{w}(\tilde{z})$ following the original framework (gorinova2020automatic, ). In practice, we can choose any base family $q_{w}(\tilde{z})$ and still reap the benefits of partial non-centering. For more details please refer to Algorithm 2 in the Appendix B.
@@ 2.6 Forward Autoregressive Flows
Flows offer a powerful way to extend the expressiveness of variational families, thereby reducing the KL divergence between the variational approximation and the true posterior (rezende2015variational, ; kingma2016improved, ; agrawal2020advances, ; agrawal2024disentangling, ). In flows, a simple base distribution is gradually transformed into a more complex one through a sequence of invertible mappings. One such transformation is the forward autoregressive flow.
@@ 3 Model-Informed Flows: Bridging VIP and Autoregressive Flows
We return to our central question: Can a flow-based distribution encapsulate the benefits of VIP? The VIP process, when applied in the variational space, involves two key stages: first, sampling an auxiliary variable $\tilde{z}$ from a base variational distribution $q_{w}(\tilde{z})$, and second, transforming $\tilde{z}$ into the latent variable $z$ using the VIP transformation, $T_{\mathrm{VIP}}$.
The original VIP framework (gorinova2020automatic, ) primarily considers $q_{w}(\tilde{z})$ to be a mean-field Gaussian (i.e., with a diagonal covariance matrix). Our investigation extends this by considering $q_{w}(\tilde{z})$ as a full-rank Gaussian.
@@
(see Appendix C in Appendix C). Thus, a full-rank Gaussian distribution can be viewed as a specific instance of an affine FAF.
@@
If we attempt to cast this composite transformation into the standard affine FAF form from Equation 2.12 the terms would need to be
$m_{i}(z_{1:i-1})$ =?${{=}}f_{i}(\pi(z_{i}))+g_{i}(\pi(z_{i}))^{1-\lambda_{i}}\Bigl{(}\mu_{i}+\sum_{j=1}^{i-1}L_{ij}\epsilon_{j}-\lambda_{i}f_{i}(\pi(z_{i}))\Bigr{)},$ (3.4)
$\log s_{i}(z_{1:i-1})$ =?${{=}}\log g_{i}(\pi(z_{i}))^{1-\lambda_{i}}+\log L_{ii}.$ (3.5)
In principle, sufficiently powerful networks in the FAF could represent these forms. However, they could not be captured in an affine FAF due to dependence on $\epsilon$ and prior functions as well as the quadratic form in Equation 3.4. These nonlinearities may be difficult to capture even with an FAF that uses, e.g., multi-layer perceptrons. To remedy this, we introduce a generalized flow.
@@
theoremvipdesignarbitrary Let $T=T_{\mathrm{VIP}}\circ T_{A}$ where $T_{\mathrm{VIP}}$ is the VIP transformation (Definition 1) and $T_{A}$ is the affine transformation from full-rank Gaussian. If $f_{i}$ and $\log g_{i}$ in the hierarchical Bayesian model (Equation 2.2) are arbitrary continuous functions, and the parent nodes $\pi(z_{i})$ are a subset of the preceding variables $z_{1:i-1}$ (respecting the causal dependency structure of the model), then $T$ can be represented as an affine generalized forward autoregressive flow with
@@
Definition 3 therefore demonstrates that GFAF, by leveraging its novel translation term $t_{i}$ (which depends on past noise $\epsilon_{1:i-1}$) and by incorporating the model’s prior functions, can replicate the full-rank VIP mechanism.
@@ 3.1 Model-Informed Flows
1:$\epsilon=(\epsilon_{1},\cdots,\epsilon_{D})$ , prior functions $f_{1},\cdots,f_{D}$ and $g_{1},\cdots,g_{D}$ (or their logs for scale)
@@
3:     $u_{i}\leftarrow[z_{1:i-1}{\color[rgb]{0,0,1},f_{i}(\pi(z_{i})),\log g_{i}(\pi(z_{i}))}]$triangleright Construct input incorporating prior information
4:     $m_{i}\leftarrow\text{NN}_{m}(u_{i})$triangleright Neural network for the shift function
5:     $\log s_{i}\leftarrow\text{NN}_{s}(u_{i})$triangleright Neural network for the log-scale function
6:     $t_{i}\leftarrow\text{NN}_{t}([u_{i},\epsilon_{1:i-1}])$triangleright Neural network for the translation term with $\epsilon_{1:i-1}$ inputs
7:     $z_{i}\leftarrow m_{i}+\exp(\log s_{i})\cdot\Bigl{(}\epsilon_{i}{\color[rgb]{0,0,1}-t_{i}}\Bigr{)}$triangleright Generalized autoregressive transformation
@@
Algorithm 1 presents the pseudocode for implementing MIF, highlighting in blue the components that extend a standard forward autoregressive flow; omitting these reduces MIF to such a standard FAF. The key architectural differences are, first, MIF employs conditioning networks not only for the shift $m_{i}$ and log-scale $\log s_{i}$ but also for the translation term $t_{i}$, with the latter explicitly taking previous noise variables $\epsilon_{1:i-1}$ as inputs, as motivated by Definition 3. Second, MIF is designed to incorporate the model’s prior functions $f_{i}$ and $\log g_{i}$ as additional inputs to these conditioning networks, providing them with valuable, explicit structural guidance from the target model. Third, the generation process in MIF (the loop over $i$) must adhere to a pre-defined topological order of the latent variables consistent with the hierarchical Bayesian model.
@@ 3.2 Incorporating Prior Information through $f_{i}$ and $\log g_{i}$
Our preliminary experiments revealed that omitting explicit prior function inputs ($f_{i},\log g_{i}$) from MIF often minimally impacted performance. This occurs when these priors are affine, as the target generalized forward autoregressive flow (GFAF) representing full-rank VIP (per Definition 3) then simplifies to an affine GFAF. An MIF, even without these explicit prior inputs, can learn to reproduce this affine GFAF if its own conditioning functions for $m_{i},\log s_{i},$ and $t_{i}$ possess at least affine capacity. This is shown in the following corollary.
@@
corollaryvipdesignaffine Let $T=T_{\mathrm{VIP}}\circ T_{A}$ where $T_{\mathrm{VIP}}$ is the VIP transformation (see Definition 1) and $T_{A}$ is the affine transformation from full-rank Gaussian. If $f_{i}$ and $\log g_{i}$ in the hierarchical Bayesian model (see Equation 2.2) are affine functions, then $T$ can be represented as an affine generalized forward autoregressive flow.

### v2

# Model-Informed Flows for Bayesian Inference

@@ Abstract
Variational inference often struggles with the posterior geometry exhibited by complex hierarchical Bayesian models. Recent advances in flow‐based variational families and Variationally Inferred Parameters (VIP) each address aspects of this challenge, but their formal relationship is unexplored. Here, we prove that the combination of VIP and a full-rank Gaussian can be represented exactly as a forward autoregressive flow augmented with a translation term and input from the model’s prior. Guided by this theoretical insight, we introduce the Model‐Informed Flow (MIF) architecture, which adds the necessary translation mechanism, prior information, and hierarchical ordering. Empirically, MIF delivers tighter posterior approximations and matches or exceeds state‐of‐the‐art performance across a suite of hierarchical and non‐hierarchical benchmarks.
@@ 1 Introduction
Inference remains challenging for many complex Bayesian models. Variational Inference (VI) casts posterior approximation as minimization of the Kullback-Leibler divergence between a tractable approximating distribution and the true posterior (blei2017variational; jordan1999introduction; hoffman2013stochastic). Recent advances have focused on increasing the expressiveness of variational families. Flow-based models, which apply a sequence of invertible transformations to a simple base distribution, have shown particular promise (papamakarios2017masked; rezende2015variational; kingma2016improved).
Variationally Inferred Parameters (VIP) (gorinova2020automatic) address a key challenge in VI: The posteriors of hierarchical Bayesian models often exhibit pronounced curvature or “funnel-like” geometry (betancourt2015hamiltonian; ingraham2017variational) that standard families like Gaussians cannot capture. VIP builds on insights from non-centered parameterization (NCP) in Markov chain Monte Carlo (MCMC) (papaspiliopoulos2007general), adaptively learning the optimal degree of non-centering for each latent variable during the VI optimization process.
Our preliminary investigations reveal a striking observation: When VIP is combined with full-rank Gaussian variational families, it achieves performance comparable to state-of-the-art methods from recent evaluation (blessing2024beyond) (See Tables 1 and  3). Since VIP and flow-based models both excel at navigating posterior geometry, might they be manifestations of a common principle? In particular, can a flow-based distribution encapsulate the benefits of VIP?
@@ 2.2 Hierarchical Bayesian Models
Hierarchical Bayesian models are common in domains such as ecology (cressie2009accounting; wikle2003hierarchical; royle2008hierarchical) and epidemiology (dominici2000combining; richardson2003bayesian; lawson2018bayesian). Following Gorinova et al. gorinova2020automatic, we consider a hierarchical model in which each latent variable $z_{i}$ is conditionally Gaussian with a mean and variance that are arbitrary function of parent variables, while the observed data $x_{j}$ is generated from an arbitrary likelihood $p(x_{j}|\pi(x_{j}))$. In particular, the likelihood for $x_{j}$ may depend on one or more latent variables through its parent set $\pi(x_{j})$. Formally, write
@@
where $\pi(\cdot)$ denotes the set of parent variables. Throughout, $\pi(\cdot)$ may include both observed and latent variables; that is, a variable can be conditioned on any subset of the remaining variables permitted by the model graph. In practice, the functions $f_{i}$ and $g_{i}$ (which return the mean and standard deviation of $z_{i}$, respectively) often arise directly from a probabilistic programming system, encoding how the mean and variance of each latent variable depend on its parent variables.
@@
Example.
As a concrete instance of  Equation˜2.2, consider the model
$z_{1}$ $\sim\mathcal{N}(0,1),$ $\qquad z_{2}$ $\sim\mathcal{N}\!\bigl(\alpha z_{1},\,\exp(\beta z_{1})\bigr),$
$x_{1}\mid z_{1},z_{2}$ $\sim\mathrm{Poisson}\!\left(\exp\!\bigl(z_{1}+\tfrac{1}{2}z_{2}\bigr)\right),$ $\qquad x_{2}\mid z_{2},x_{1}$ $\sim\mathrm{Binomial}\!\left(n=x_{1},\,p=\sigma(z_{2})\right),$
where $\sigma(u)=\frac{1}{1+\exp(-u)}$. To cast this as an instance of Equation˜2.2, define the parent sets
$\pi(z_{1})=\emptyset,\quad\pi(z_{2})=\{z_{1}\},\quad\pi(x_{1})=\{z_{1},z_{2}\},\quad\pi(x_{2})=\{z_{2},x_{1}\},$
and the functions
$f_{1}(\emptyset)=0,\quad g_{1}(\emptyset)=1,\qquad f_{2}(z_{1})=\alpha z_{1},\quad g_{2}(z_{1})=\exp(\beta z_{1}).$
In Equation˜2.2, we make no assumptions about the distributional form of the observed variables $x_{i}$, so $p(x_{1}|\pi(x_{1}))=\mathrm{Poisson}\!\left(\exp\!\bigl(z_{1}+\tfrac{1}{2}z_{2}\bigr)\right)$ and $p(x_{2}|\pi(x_{2}))=\mathrm{Binomial}\!\left(n=x_{1},\,p=\sigma(z_{2})\right)$ can be arbitrary.
@@ 2.3 Non-Centered Parameterization
Hierarchical models often exhibit difficult curvature and “funnel-like“ geometries. These are a challenge for both sampling and variational inference approaches (betancourt2015hamiltonian; ingraham2017variational; hoffman2019neutra). A common strategy in MCMC to alleviate these issues is the non-centered parameterization (NCP) (papaspiliopoulos2007general). To illustrate, consider two latent variables $z_{1}$ and $z_{2}$ having prior of
@@
In the standard (centered) parameterization, the goal is to approximate $p(z|x)$. However, the nonlinear dependence of $z_{2}$ on $z_{1}$ can yield a sharply funnel-shaped joint posterior. As illustrated in Figure˜1 (left panel), a full-rank Gaussian struggles to capture this funnel structure.
@@ 2.4 Variationally Inferred Parameters
then recovers $z_{i}$ via VIP transformation, $T_{\mathrm{VIP}}$, as in Definition˜1. Note that $\lambda_{i}$ interpolates between the centered ($\lambda_{i}=1$) and fully non-centered ($\lambda_{i}=0$) extremes.
@@
Substituting $T_{\mathrm{VIP}}$ into the joint distribution (Equation˜2.3) yields a partially non-centered representation
@@
[…]
@@ 2.5 Equivalence of VIP in Model vs Variational Spaces
[…]
@@
Since the KL divergence is invariant under such invertible transformations (cover1991elements), we have that
@@
[…]
@@ 2.6 Forward Autoregressive Flows
Flows offer a powerful way to extend the expressiveness of variational families, thereby reducing the KL divergence between the variational approximation and the true posterior (rezende2015variational; kingma2016improved; agrawal2020advances; agrawal2024disentangling). In flows, a simple base distribution is gradually transformed into a more complex one through a sequence of invertible mappings. One such transformation is the forward autoregressive flow.
@@ 3 Model-Informed Flows: Bridging VIP and Autoregressive Flows
[…]
@@
(see Appendix˜C in Appendix˜C). Thus, a full-rank Gaussian distribution can be viewed as a specific instance of an affine FAF.
@@
If we attempt to cast this composite transformation into the standard affine FAF form from Equation˜2.12 the terms would need to be
$m_{i}(z_{1:i-1})$ $\stackrel{{\scriptstyle?}}{{=}}f_{i}(\pi(z_{i}))+g_{i}(\pi(z_{i}))^{1-\lambda_{i}}\Bigl(\mu_{i}+\sum_{j=1}^{i-1}L_{ij}\epsilon_{j}-\lambda_{i}f_{i}(\pi(z_{i}))\Bigr),$ (3.4)
$\log s_{i}(z_{1:i-1})$ $\stackrel{{\scriptstyle?}}{{=}}\log g_{i}(\pi(z_{i}))^{1-\lambda_{i}}+\log L_{ii}.$ (3.5)
In principle, sufficiently powerful networks in the FAF could represent these forms. However, they could not be captured in an affine FAF due to dependence on $\epsilon$ and prior functions. In addition, Equation˜3.4 contains the product $g_{i}(\pi(z_{i}))\,f_{i}(\pi(z_{i}))$, which induces quadratic dependence on $z$ that an affine mapping cannot represent. These nonlinearities may be difficult to capture even with an FAF that uses, e.g., multi-layer perceptrons. To remedy this, we introduce a generalized flow.
@@
theoremvipdesignarbitrary Let $T=T_{\mathrm{VIP}}\circ T_{A}$ where $T_{\mathrm{VIP}}$ is the VIP transformation (Definition˜1) and $T_{A}$ is the affine transformation from full-rank Gaussian. If $f_{i}$ and $\log g_{i}$ in the hierarchical Bayesian model (Equation˜2.2) are arbitrary continuous functions, and the parent nodes $\pi(z_{i})$ are a subset of the preceding variables $z_{1:i-1}$ (respecting the causal dependency structure of the model), then $T$ can be represented as an affine generalized forward autoregressive flow with
@@
Definition˜3 therefore demonstrates that GFAF, by leveraging its novel translation term $t_{i}$ (which depends on past noise $\epsilon_{1:i-1}$) and by incorporating the model’s prior functions, can replicate the full-rank VIP mechanism.
@@ 3.1 Model-Informed Flows
[…]
@@
3:  $u_{i}\leftarrow[z_{1:i-1}{\color[rgb]{0,0,1},f_{i}(\pi(z_{i})),\log g_{i}(\pi(z_{i}))}]$ $\triangleright$ Construct input incorporating prior information
4:  $m_{i}\leftarrow\text{NN}_{m}(u_{i})$ $\triangleright$ Neural network for the shift function
5:  $\log s_{i}\leftarrow\text{NN}_{s}(u_{i})$ $\triangleright$ Neural network for the log-scale function
6:  $t_{i}\leftarrow\text{NN}_{t}([u_{i},\epsilon_{1:i-1}])$ $\triangleright$ Neural network for the translation term with $\epsilon_{1:i-1}$ inputs
7:  $z_{i}\leftarrow m_{i}+\exp(\log s_{i})\cdot\Bigl(\epsilon_{i}{\color[rgb]{0,0,1}-t_{i}}\Bigr)$ $\triangleright$ Generalized autoregressive transformation
@@
Algorithm˜1 presents the pseudocode for implementing MIF, highlighting in blue the components that extend a standard forward autoregressive flow; omitting these reduces MIF to such a standard FAF. The key architectural differences are, first, MIF employs conditioning networks not only for the shift $m_{i}$ and log-scale $\log s_{i}$ but also for the translation term $t_{i}$, with the latter explicitly taking previous noise variables $\epsilon_{1:i-1}$ as inputs, as motivated by Definition˜3. Second, MIF is designed to incorporate the model’s prior functions $f_{i}$ and $\log g_{i}$ as additional inputs to these conditioning networks, providing them with valuable, explicit structural guidance from the target model. Third, the generation process in MIF (the loop over $i$) must adhere to a pre-defined topological order of the latent variables consistent with the hierarchical Bayesian model.
@@ 3.2 Incorporating Prior Information through $f_{i}$ and $\log g_{i}$
Our preliminary experiments revealed that omitting explicit prior function inputs ($f_{i},\log g_{i}$) from MIF often minimally impacted performance. This occurs when these priors are affine, as the target generalized forward autoregressive flow (GFAF) representing full-rank VIP (per Definition˜3) then simplifies to an affine GFAF. An MIF, even without these explicit prior inputs, can learn to reproduce this affine GFAF if its own conditioning functions for $m_{i},\log s_{i},$ and $t_{i}$ possess at least affine capacity. This is shown in the following corollary.
@@
corollaryvipdesignaffine Let $T=T_{\mathrm{VIP}}\circ T_{A}$ where $T_{\mathrm{VIP}}$ is the VIP transformation (see Definition˜1) and $T_{A}$ is the affine transformation from full-rank Gaussian. If $f_{i}$ and $\log g_{i}$ in the hierarchical Bayesian model (see Equation˜2.2) are affine functions, then $T$ can be represented as an affine generalized forward autoregressive flow.